Can you help me solve this equation -x+y=-13 3x-y=19

Answer by stanbon(75887)

Can you help me solve this system of equations::
-x+y=-13
3x-y=19
----
Add and solve for "y"::
2x + 0 = 6
---
x = 3
-------
Solve for "y":
-x + y = -13
-3 + y = -13
y = -10
--------------
Cheers,
Stan H.
-------------